# Continuous Domain Generalization

### v1

@@ 4.2 Neural Domain Transport Operator under Structured Constraints
The neural transport operator $\mathcal{T}$ is said to satisfy an algebraic structure if the following properties hold: Closure: $\mathcal{T}(\theta(z_{i}),z_{i},z_{j})\in\Theta$; Identity: $\mathcal{T}(\theta(z_{i}),z_{i},z_{i})=\theta(z_{i})$; Associativity: $\mathcal{T}(\mathcal{T}(\theta(z_{i}),z_{i},z_{j}),z_{j},z_{k})=\mathcal{T}(\theta(z_{i}),z_{i},z_{k})$; Invertibility: $\mathcal{T}^{-1}\mathcal{T}(\theta(z_{i}),z_{i},z_{j})=\theta(z_{i})$. Here, closure ensures that transported parameters always remain in the valid parameter space. Identity ensures that transporting to the same descriptor leaves parameters unchanged. Associativity ensures that sequential transports match the direct one. Invertibility ensures that each transport can be exactly reversed.
The geometric structure ensures that $\theta(z)$ forms a smooth manifold, as established in Theorem 1. The algebraic structure defined via Group-like axioms promotes Equivariance. Equivariance ensures that for any two domains $\mathcal{D}_{i},\mathcal{D}_{j}$, the operator yields a predictor whose behavior aligns with the semantic shift from $P(Y|X;z_{i})$ to $P(Y|X;z_{j})$. Without equivariance, the learned operator may produce parameter trajectories that are inconsistent with the underlying evolution of the data distribution.
@@
The geometric and algebraic structure jointly imply that the parameter family =$\mathcal{M}=\{\theta(z)\mid z\in\mathcal{Z}\}$ forms Lie Group—i.e., a smooth manifold with a compatible group operation. The transport operator induces parameter Lie group transitions along directions specified by domain descriptors. This motivates us to propose the Neural Lie Transport Operator (NeuralLTO), a learnable operator grounded in Lie theory and parameterized by a neural network. NeuralLTO characterizes the local variation of $\theta(z)$ using the Lie algebra $\mathfrak{g}$ generated from the domain descriptor. The Lie algebra $\mathfrak{g}$ serves as a tangent space that captures the differential structure of the parameter manifold. Given a source descriptor $z_{i}$ and corresponding parameter $\theta(z_{i})$, we generate a Lie algebra element $V(z_{i})\in\mathfrak{g}$ using a neural network. For a target descriptor $z_{j}$, we apply the offset $(z_{j}-z_{i})$ to $V(z_{i})$ to transport $\theta(z_{i})$ to $\theta(z_{j})$:
@@
In practice, the parameter vector $\theta(z)\in\mathbb{R}^{D}$ is high-dimensional. We first encode it into a compact latent representation $e(z)=f_{e}(\theta(z))\in\mathbb{R}^{m}$ using an encoder $f_{e}$. We then perform cascaded transport in the latent space. Specifically, we define a set of $d$ neural Lie algebra fields $\{V^{1},\dots,V^{d}\}$, where each $V^{k}\in\mathbb{R}^{m\times m}$ is produced by a field network $f_{v}^{k}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{m\times m}$ conditioned on the source descriptor $z_{i}$. Each $f_{v}^{k}$ can be instantiated with a generic deep neural network. Given a descriptor shift $(z_{j}-z_{i})\in\mathbb{R}^{d}$, we compute the cumulative transformation as:
@@ 4.3 Handling Imperfect Descriptors Space
In the real world, the descriptor space $\mathcal{Z}$ may be imperfect: it can be noisy, redundant, or incomplete. These imperfections lead to two principal challenges: (1) Degeneracy: Some directions in $\mathcal{Z}$ fail to affect $\theta(z)$, or redundantly encode variations. (2) Incompleteness: The observed descriptors may partially capture the true degrees of freedom underlying domain shifts, leaving latent factors unmodeled.
@@
When the descriptor $z$ is incomplete, learning a globally consistent transport function becomes underdetermined. We adopt a localized modeling strategy grounded in the theory of differential geometry: the parameter manifold is represented as an atlas, i.e., a collection of overlapping local charts. For each descriptor $z_{i}$, we define a neighborhood:
@@
We optimize NeuralLTO by supervising the transport operator $\mathcal{T}$ over local neighborhoods. Specifically, each descriptor $z_{i}$ is paired with its neighbors $z_{j}\in\mathcal{N}(z_{i})$, and $\mathcal{T}$ is trained to map parameters across these pairs. At test time, we infer $\theta(z_{s})$ by transporting from the nearest training descriptor. Detailed procedures for both training and inference are summarized in Appendix B. We also provide a detailed model complexity analysis in Appendix A.4.
@@ 5 Experiment
In this section, we evaluate the effectiveness of the proposed framework across diverse continuous domain generalization tasks. Our experimental study is designed to answer the following key questions: 1) Can the framework effectively generalize predictive models across continuously evolving domains? 2) Can the framework faithfully recover the underlying parameter manifold shaped by domain variation? 3) Can the framework robustly handle descriptor imperfections? More detailed results (i.e., dataset details, baseline details, model and hyperparameter configurations, ablation study, scalability analysis, sensitivity analysis, and convergence analysis) are demonstrated in Appendix A.
Synthetic Datasets. Two synthetic datasets are employed to simulate continuous domain shifts under interpretable variations. In the 2-Moons dataset, each domain is generated by applying scaling and rotation to the base moon shape. The descriptor $z=[z^{1},z^{2}]\in\mathbb{R}^{2}$ encodes the scale factor and rotation angle, respectively. We train on 50 randomly sampled domains (see Fig. 2), and test on 150 others plus a fixed mesh grid of descriptor points (see Fig. 5). In the MNIST dataset, each domain is constructed by applying rotation, color shift (red to blue), and occlusion to digits. The descriptor $z=[z^{1},z^{2},z^{3}]\in\mathbb{R}^{3}$ encodes the intensity of each transformation. We use 50 randomly selected domains for training and another 50 for testing. Visual illustrations are shown in Fig. 6.
Real-world Datasets. We further evaluate on a collection of diverse, multimodal real-world datasets: fMoW, ArXiv, Yearbook for classification, and Traffic for regression. For each dataset, we collect auxiliary information from metadata or context to construct descriptors. Full details of the dataset and descriptor construction are provided in Appendix A.1.
@@
ERM 34.7 $\pm$ 0.2 31.8 $\pm$ 0.9 37.1 $\pm$ 0.4 35.6 $\pm$ 0.1 8.6 $\pm$ 1.0 16.4 $\pm$ 0.1
IRM [1] 34.4 $\pm$ 0.2 33.0 $\pm$ 0.8 48.3 $\pm$ 3.9 37.4 $\pm$ 1.0 8.3 $\pm$ 0.5 16.6 $\pm$ 0.1
V-REx [26] 34.9 $\pm$ 0.1 32.2 $\pm$ 1.4 38.6 $\pm$ 0.7 37.3 $\pm$ 0.7 8.9 $\pm$ 0.5 20.9 $\pm$ 0.6
GroupDRO [46] 34.5 $\pm$ 0.1 37.6 $\pm$ 1.0 38.1 $\pm$ 1.0 35.6 $\pm$ 0.1 8.0 $\pm$ 0.4 16.2 $\pm$ 0.1
Mixup [54] 34.9 $\pm$ 0.1 34.0 $\pm$ 0.9 37.1 $\pm$ 1.0 35.5 $\pm$ 0.2 7.5 $\pm$ 0.5 16.1 $\pm$ 0.1
DANN [17] 35.1 $\pm$ 0.4 34.7 $\pm$ 0.6 38.7 $\pm$ 1.8 36.5 $\pm$ 0.2 8.9 $\pm$ 1.4 18.1 $\pm$ 0.2
MLDG [27] 34.6 $\pm$ 0.2 85.1 $\pm$ 2.5 39.2 $\pm$ 0.4 35.8 $\pm$ 0.2 7.7 $\pm$ 0.5 16.9 $\pm$ 0.1
CDANN [32] 35.0 $\pm$ 0.2 36.4 $\pm$ 0.8 38.2 $\pm$ 0.3 36.2 $\pm$ 0.2 8.7 $\pm$ 0.4 17.3 $\pm$ 0.2
URM [25] 34.7 $\pm$ 0.1 31.8 $\pm$ 1.3 38.7 $\pm$ 0.6 35.5 $\pm$ 0.4 8.0 $\pm$ 0.3 16.2 $\pm$ 0.2
@@ 5.1 Quantitative analysis
We present the performance of our proposed method against baseline methods, highlighting results from Table 1. Our method demonstrates strong generalizability across a wide range of continuous domains, with particularly large improvements on 2-Moons (3.2% vs. 13.1%) and MNIST (9.5% vs. 26.3%), indicating its effectiveness in modeling structured domain variation. Several key observations emerge: (1) Descriptor-aware methods perform effectively. However, naïvely incorporating domain descriptors (e.g., ERM-D) fails to consistently improve performance, as it may introduce spurious descriptor biases that hinder generalization. (2) Temporal methods are insufficient for general continuous domains. Models designed for temporal domain generalization (i.e., TKNets, DRAIN, Koodos) rely on domain sequence and fail when the descriptor space lacks such an ordering structure. (3) Continuous domain generalization is intrinsically challenging. Even among descriptor-aware methods, the performance of baselines on synthetic datasets is unsatisfactory. No existing method achieves reasonable accuracy. This highlights that prior methods largely overlook the underlying structure of domain variation, which is essential for modeling continuous generalization.
Figure 3: Visualization of the learned parameter manifold and the corresponding generalization behavior. Left: PCA projection of predicted parameters $\hat{\theta}(z)$ over the entire descriptor space. Right: Visualization of decision boundaries and data samples along selected direction.
@@ 5.2 Qualitative analysis
To understand how domain descriptors influence model behavior, we visualize the decision boundaries on the 2-Moons dataset. In Fig. 2, we randomly sample four unseen test domains (marked as blue crosses in the descriptor space) and visualize the classification boundaries of three baselines on the right panel. Each row corresponds to a method, and each column to a test domain. ERM-D embeds descriptors directly into the model input, leading to entangled representations. This often induces spurious correlations and overfitting, resulting in highly irregular and fragmented decision boundaries. NDA relies solely on chance overlap between training and test domains, lacking any generalization mechanism. As a result, it produces unstable decision boundaries with no guarantee. CIDA performs adversarial alignment based on descriptor distance and partially captures scaling, seen as ring expansion. However, it fails to capture rotation, as the underlying metric lacks structural fidelity. This limitation reflects the inadequacy of distance-based alignment in preserving domain geometry, particularly when transformations are structurally non-isometric.
Fig. 3 presents a visual analysis of the learned parameter manifold of our model in the 2-Moons dataset. Specifically, we train the model using the given training domains, and then densely generalize $\theta(z)$ across the full descriptor domain. On the left, we apply PCA to the predicted model parameters and project them over the 2D descriptor plane. The resulting surface is continuous, semantically aligned, and structurally coherent. On the right, we visualize decision boundaries along three representative directions: scaling-only (black), rotation-only (red), and mixed (blue). For each, the models exhibit clear and concise boundaries that adapt appropriately to the changes in the data distribution, firmly demonstrating that our method captures the structural geometry of domain variation and enables interpretable generalization across the descriptor space.
@@ 5.3 Imperfections analysis
We design three controlled experiments to simulate the common imperfections of the domain descriptor: noise, redundancy, and incompleteness, as challenges discussed in Section 4.3. To assess the noise, we augment $z$ with randomly sampled noise dimensions, forming descriptors with uninformative components. To assess redundancy, we apply redundant projections to the original descriptor to obtain a higher-dimensional, over-parameterized version. This setup mimics real-world cases where descriptors are overly entangled. To study incompleteness, we simulate missing information by randomly dropping $n$ dimensions from the redundant descriptor. As $n$ increases, the retained dimensions provide less faithful representation of the underlying domain variation. An overview of the construction is illustrated in the top row of Fig. 4.
@@
The bottom row of Fig. 4 shows the error rates under increasing levels of descriptor imperfection across three settings. In the Noisy setting (left), we gradually append up to five random noise dimensions to the original descriptor. As noise increases, the performance of the model without gating rapidly deteriorates, reaching over 30% error at the highest noise level. In contrast, the model equipped with our gating mechanism remains significantly more stable. This indicates that although the noise dimensions are uninformative, the remaining structure in the descriptor space still guides transport reasonably well, provided that irrelevant dimensions are actively suppressed. In the Redundant setting (middle), we expand the descriptor by applying random linear projections to the original descriptor, introducing redundancy without discarding any information. Both models maintain reasonably low error rates (under 10%), since the underlying domain variation remains fully preserved. However, the model equipped with gating achieves slightly better and more stable performance. This suggests that while redundancy itself is not harmful, the gating mechanism helps by softly recovering the original independent factors of variation, effectively regularizing the projection space and improving robustness. In the Incomplete setting (right), we construct an 8-dimensional descriptor space embedding two latent factors of variation, and simulate increasing incompleteness by progressively removing 1 to 7 dimensions. The model without charting (No-Chart) fails once key structural information is removed. In contrast, our chart-based variant exhibits a controlled performance decline. This validates the benefit of local manifold modeling: by restricting transport computation to descriptor neighborhoods, the chart-based approach preserves continuity even when global descriptor integrity is compromised.

### v2

@@ 4.2 Neural Domain Transport Operator under Structural Constraints
The neural transport operator $\mathcal{T}$ is said to satisfy algebraic structure if the following properties hold:
  • Closure: $\mathcal{T}(\theta(z_{i}),z_{i},z_{j})\in\Theta$, ensuring transported parameters remain within the valid space.
  • Identity: $\mathcal{T}(\theta(z_{i}),z_{i},z_{i})=\theta(z_{i})$, ensuring that self-transport leaves parameters unchanged.
  • Associativity: $\mathcal{T}(\mathcal{T}(\theta(z_{i}),z_{i},z_{j}),z_{j},z_{k})=\mathcal{T}(\theta(z_{i}),z_{i},z_{k})$, ensuring that sequential transports are equivalent to direct ones.
  • Invertibility: $\mathcal{T}^{-1}\!\mathcal{T}(\theta(z_{i}),z_{i},z_{j})=\theta(z_{i})$, ensuring that each transport can be exactly reversed.
The geometric structure ensures that $\theta(z)$ forms a smooth manifold, as established in Theorem 1. The algebraic structure defined via Group-like axioms imposes Equivariance. Equivariance characterizes a symmetry relation between transformations in the input and output spaces. In this context, it ensures that applying a transformation in the descriptor space and then mapping to parameters yields the same result as transforming parameters directly. Without equivariance, the learned operator may produce parameter trajectories that deviate from the true evolution of the underlying data distribution.
@@
The geometric and algebraic structure jointly imply that the parameter family $\mathcal{M}=\{\theta(z)\mid z\in\mathcal{Z}\}$ forms Lie Group—i.e., a smooth manifold with a compatible group operation. The transport operator induces parameter Lie Group transitions along directions specified by domain descriptors. This motivates us to propose the Neural Lie Transport Operator (NeuralLio), a learnable operator grounded in Lie theory and parameterized by a neural network. NeuralLio characterizes the local variation of $\theta(z)$ using the Lie algebra $\mathfrak{g}$ generated from the domain descriptor. The Lie algebra $\mathfrak{g}$ serves as a tangent space that captures the differential structure of the parameter manifold. Given a source descriptor $z_{i}$ and corresponding parameter $\theta(z_{i})$, we generate a Lie algebra element $V(z_{i})\in\mathfrak{g}$ using a neural network. For a target descriptor $z_{j}$, we apply the offset $(z_{j}-z_{i})$ to $V(z_{i})$ to transport $\theta(z_{i})$ to $\theta(z_{j})$:
@@
In practice, the parameter vector $\theta(z)\in\mathbb{R}^{D}$ is high-dimensional. We first encode it into a compact latent representation $e(z)=f_{e}(\theta(z))\in\mathbb{R}^{m}$ using an encoder $f_{e}$. We then perform cascaded transport in the latent space. Specifically, we define a set of $d$ neural Lie algebra fields $\{V^{1},\dots,V^{d}\}$, where each $V^{k}\in\mathbb{R}^{m\times m}$ is produced by a field network $f_{v}^{k}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{m\times m}$ conditioned on the source descriptor $z_{i}$. Each $f_{v}^{k}$ can be instantiated with a generic deep neural network. Given a descriptor shift $(z_{j}-z_{i})$, we compute the cumulative transformation as:
@@ 4.3 Handling Imperfect Descriptors Space
Theorem 1 discusses a theoretically grounded descriptor space. In the real world, however, the descriptors accessible for model training are observable projections of this theoretical space. They may contain noise, redundancy, or missing factors, leading to imperfect representations. These imperfections introduce two challenges: (1) Degeneracy: Some directions in the observed $\mathcal{Z}$ fail to affect $\theta(z)$ or redundantly encode variations. (2) Incompleteness: The observed descriptors partially capture the degrees of freedom governing domain shifts, leaving some latent factors unmodeled.
@@
When the space $\mathcal{Z}$ is incomplete, learning a globally consistent transport function becomes underdetermined. We adopt a localized modeling strategy grounded in the theory of differential geometry: the parameter manifold is represented as an atlas, i.e., a collection of overlapping local charts. For each descriptor $z_{i}$, we define its neighborhood as:
@@
NeuralLio is optimized by supervising the transport operator $\mathcal{T}$ over local neighborhoods in the descriptor space. For each training descriptor $z_{i}$, we sample nearby descriptors $z_{j}\in\mathcal{N}(z_{i})$ and train $\mathcal{T}$ to transport model parameters from $z_{i}$ to $z_{j}$ so that the resulting predictions and latent representations remain consistent. During inference, parameters for an unseen domain $z_{s}$ are inferred by transporting from the nearest training descriptor using the learned operator. Detailed procedures for both training and inference are summarized in Appendix B. We also provide a detailed model complexity analysis in Appendix A.4.
@@ 5 Experiment
In this section, we evaluate the effectiveness of the proposed framework across diverse continuous domain generalization tasks. Our experimental study is designed to answer the following key questions: 1) Can the framework effectively generalize predictive models across continuously evolving domains? 2) Can the framework faithfully recover the underlying parameter manifold shaped by domain variation? 3) Can the framework robustly handle descriptor imperfections? 4) Can the learned transport operator exhibit the desired structural properties? More detailed results (i.e., dataset details, baseline details, model and hyperparameter configurations, ablation study, scalability analysis, sensitivity analysis, and convergence analysis) are demonstrated in Appendix A.
Synthetic Datasets. Two synthetic datasets are employed to simulate continuous domain shifts under interpretable variations. In the 2-Moons dataset, each domain is generated by applying scaling and rotation to the base moon shape. The descriptor $z=[z^{1},z^{2}]\in\mathbb{R}^{2}$ encodes the scale factor and rotation angle, respectively. We train the model on 50 randomly sampled domains, and evaluate it on 150 additional randomly sampled domains, together with extra test domains uniformly sampled over a mesh grid in the descriptor space (see Fig. 5). In the MNIST dataset, each domain is constructed by applying rotation, color shift (red to blue), and occlusion to digits. The descriptor $z=[z^{1},z^{2},z^{3}]\in\mathbb{R}^{3}$ encodes the intensity of each transformation. We use 50 randomly selected domains for training and another 50 for testing. Visual illustrations are shown in Fig. 6.
Real-world Datasets. We further evaluate our framework on a diverse collection of multimodal real-world datasets: fMoW, ArXiv, and Yearbook for classification, and Traffic for regression. For each dataset, we construct continuous descriptors using auxiliary information derived from publicly available domain metadata or contextual variables. Comprehensive details of dataset composition, processing, and descriptor construction are provided in Appendix A.1.
@@
ERM 34.7 $\pm$ 0.2 31.8 $\pm$ 0.9 27.7 $\pm$ 1.6 35.6 $\pm$ 0.1 8.6 $\pm$ 1.0 16.4 $\pm$ 0.1
IRM [1] 34.4 $\pm$ 0.2 33.0 $\pm$ 0.8 41.5 $\pm$ 2.8 37.4 $\pm$ 1.0 8.3 $\pm$ 0.5 16.6 $\pm$ 0.1
V-REx [26] 34.9 $\pm$ 0.1 32.2 $\pm$ 1.4 32.1 $\pm$ 3.6 37.3 $\pm$ 0.7 8.9 $\pm$ 0.5 20.9 $\pm$ 0.6
GroupDRO [46] 34.5 $\pm$ 0.1 37.6 $\pm$ 1.0 28.6 $\pm$ 1.9 35.6 $\pm$ 0.1 8.0 $\pm$ 0.4 16.2 $\pm$ 0.1
Mixup [54] 34.9 $\pm$ 0.1 34.0 $\pm$ 0.9 27.1 $\pm$ 1.5 35.5 $\pm$ 0.2 7.5 $\pm$ 0.5 16.1 $\pm$ 0.1
DANN [17] 35.1 $\pm$ 0.4 34.7 $\pm$ 0.6 26.0 $\pm$ 0.7 36.5 $\pm$ 0.2 8.9 $\pm$ 1.4 18.1 $\pm$ 0.2
MLDG [27] 34.6 $\pm$ 0.2 85.1 $\pm$ 2.5 29.2 $\pm$ 1.0 35.8 $\pm$ 0.2 7.7 $\pm$ 0.5 16.9 $\pm$ 0.1
CDANN [32] 35.0 $\pm$ 0.2 36.4 $\pm$ 0.8 27.6 $\pm$ 0.9 36.2 $\pm$ 0.2 8.7 $\pm$ 0.4 17.3 $\pm$ 0.2
URM [25] 34.7 $\pm$ 0.1 31.8 $\pm$ 1.3 26.9 $\pm$ 1.0 35.5 $\pm$ 0.4 8.0 $\pm$ 0.3 16.2 $\pm$ 0.2
@@
Figure 3: Visualization of the learned parameter manifold and the corresponding generalization behavior. Left: PCA projection of predicted parameters $\theta(z)$ over the entire descriptor space. Right: Visualization of decision boundaries and data samples along selected direction.
@@ 5.1 Quantitative Analysis
We present the performance of our proposed method against baseline methods, highlighting results from Table 1. Our method demonstrates strong generalizability across a wide range of continuous domains, with particularly large improvements on 2-Moons (3.2% vs. 13.1%) and MNIST (9.5% vs. 26.3%), indicating its effectiveness in modeling continuous domain variation. Beyond synthetic data, NeuralLio also achieves the best results on fMoW, ArXiv, Yearbook, and Traffic, indicating its robustness in handling diverse and irregular real-world domain variation. Table 1 also reveals several insights into how existing methods behave: (1) Descriptor-aware methods generally outperform descriptor-agnostic ones, yet naïvely incorporating descriptors (e.g., ERM-D) risks spurious correlations that hinder generalization. (2) Temporal methods are insufficient for general continuous domains. Models designed for temporal domain generalization (i.e., TKNets, DRAIN, Koodos) rely on domain sequence and fail when the descriptor space lacks such an ordering structure. (3) Continuous domain generalization is intrinsically challenging. Even among descriptor-aware methods, the performance of baselines on synthetic datasets is unsatisfactory. No existing method achieves reasonable accuracy. This highlights that prior methods largely overlook the underlying structure of domain variation, which is essential for modeling continuous generalization.
@@ 5.2 Qualitative Analysis
To further understand how domain variation influences model behavior, we visualize both baseline and learned models on the 2-Moons dataset. Fig. 2 left provides a global view of the descriptor space, where the two axes correspond to scaling (monotonic variation) and rotation (periodic variation). The background color encodes the continuous variation field across domains. All training domains are shown as black dots, with several illustrated in small inset panels. Four unseen test domains are randomly sampled for visualization and marked as blue crosses in the descriptor space. The right panel visualizes the classification boundaries of three baseline methods evaluated on the four test domains. Each row corresponds to a method, and each column to a test domain. ERM-D embeds descriptors directly into the model input, leading to entangled representations. This often induces spurious correlations and overfitting, resulting in highly irregular and fragmented decision boundaries. NDA relies solely on chance overlap between training and test domains, lacking any generalization mechanism. As a result, it produces unstable decision boundaries with no guarantee. CIDA performs adversarial alignment based on descriptor distances and partially captures the scaling effect, visible as a ring-like expansion. However, it fails to model rotational variation because its metric-driven objective lacks structural fidelity. This limitation highlights the inherent weakness of distance-based objectives in preserving domain geometry, particularly under non-isometric transformations.
Fig. 3 shows an intrinsic view of the parameter manifold learned by NeuralLio on the 2-Moons dataset. Specifically, we train the model using the given training domains, and then densely generalize $\theta(z)$ across the full descriptor space. The left panel visualizes the parameter manifold obtained by projecting all the predicted model parameters $\theta(z)$ via PCA. The resulting “twisted surface” aligns with the monotonic–periodic geometry of scaling and rotation, respectively. The right panel demonstrates that traversing along different descriptor directions corresponds to smooth, consistent transformations in decision boundaries, showing that the learned operator preserves geometric continuity and captures the co-evolution of model parameters and domain distributions. In particular, the mixed-direction case reveals the model’s ability to jointly encode multiple domain variations. In summary, the visualization results firmly demonstrate that our method captures the structural geometry of domain variation and enables interpretable generalization across the descriptor space.
@@ 5.3 Imperfections Analysis
We design three controlled experiments to simulate the common imperfections of the domain descriptor: noise, redundancy, and incompleteness, as discussed in Section 4.3. To assess the noise, we augment $z$ with randomly sampled noise dimensions, forming descriptors with uninformative components. To assess redundancy, we apply redundant projections to the original descriptor to obtain a higher-dimensional, over-parameterized version. This setup mimics real-world cases where descriptors are overly entangled. To study incompleteness, we simulate missing information by randomly dropping $n$ dimensions from the redundant descriptor. As $n$ increases, the retained dimensions provide less faithful representation of the underlying domain variation. An overview of the construction is illustrated in the top row of Fig. 4.
@@
The bottom row of Fig. 4 shows the error rates under increasing levels of descriptor imperfection across three settings. In the Noisy setting (left), we gradually append up to five random noise dimensions to the original descriptor. As noise increases, the performance of the model without gating rapidly deteriorates, reaching over 30% error at the highest noise level. In contrast, the model equipped with the gating mechanism remains significantly more stable. This indicates that although the noise dimensions are uninformative, the remaining structure in the descriptor space still guides transport reasonably well, provided that irrelevant dimensions are actively suppressed. In the Redundant setting (middle), we expand the descriptor by applying random linear projections to the original descriptor, introducing redundancy without discarding any information. Both models maintain reasonably low error rates (under 10%), since the underlying domain variation remains fully preserved. However, the model equipped with gating achieves better and more stable performance. This suggests that while redundancy itself is not harmful, the gating mechanism helps by softly recovering the original independent factors of variation, effectively regularizing the projection space and improving robustness. In the Incomplete setting (right), we construct an 8-dimensional descriptor space embedding two latent factors of variation, and simulate increasing incompleteness by progressively removing 1 to 7 dimensions. The model without charting (No-Chart) fails once key structural information is removed. In contrast, the chart-based variant exhibits a controlled performance decline. This validates the benefit of local manifold modeling, as restricting transport computation to descriptor neighborhoods allows the chart-based approach to preserve continuity even when global descriptor integrity is compromised.
5.4 Structure Property Analysis
To validate the structure properties introduced in Section 4.2, we evaluate the learned operator on the 2-Moons dataset. We measure the consistency between theoretically equivalent transformations by computing the cosine similarity of their resulting parameters across all test domains.
Table 2: Empirical verification of the structure properties of the learned operator.
Property Cosine Similarity (%)
Identity 99.9
Associativity 99.1
Invertibility 98.3
Identity. For each test descriptor $z_{i}$, we apply the self-transport $\theta(z_{i})\xrightarrow{z_{i}\to z_{i}}\theta^{\prime}(z_{i})$ and compute the cosine similarity: $\cos(\theta(z_{i}),\theta^{\prime}(z_{i}))$.
Associativity. We consider all triplet combinations $(z_{i},z_{j},z_{k})$ from the test descriptors. For each triplet, we compute two transports: $\theta(z_{i})\xrightarrow{z_{i}\to z_{j}}\theta^{\prime}(z_{j})\xrightarrow{z_{j}\to z_{k}}\theta^{\prime}(z_{k})$, and $\theta(z_{i})\xrightarrow{z_{i}\to z_{k}}\theta^{\prime\prime}(z_{k})$, then evaluate $\cos(\theta^{\prime}(z_{k}),\theta^{\prime\prime}(z_{k}))$.
Invertibility. We consider all pairs $(z_{i},z_{j})$ from the test descriptors. For each pair, we apply the round-trip transport $\theta(z_{i})\xrightarrow{z_{i}\to z_{j}}\theta^{\prime}(z_{j})\xrightarrow{z_{j}\to z_{i}}\theta^{\prime}(z_{i})$, and compute $\cos(\theta(z_{i}),\theta^{\prime}(z_{i}))$.
Closure. This property is inherently satisfied, as the operator always maps within the $\mathbb{R}^{D}$.
As shown in Table 2, the learned operator fulfills the expected properties with high precision, which empirically confirms that NeuralLio realizes the Lie-Group-based transport formulation, providing a principled foundation for structurally consistent model evolution.